OpenStudy (anonymous):

A bookshelf contains six mysteries and three biographies. Two books are selected at random without replacement. What is the probability that both books are mysteries?

OpenStudy (anonymous):

P(2 mystery) = 6C2/9C2 = 15/36 = 5/12 If I deserve medal then give me one

OpenStudy (amistre64):

there is a total of: 9! ---- ways of choosing the books 2!7! 9.8 --- = 9.4 = 36 ways total 2 If we take away the bios and see how many ways there are of picking 2 out of 6 we get: 6! 6.5 ---- = --- = 3.5 = 15 ways to get only 2 mysts. 2!4! 2 15/36 = .41666... the probablity of pick 2 mysts is then; 41' 2/3% right?